In a jib crane, the jib and tie rod are 5 m and 4 m long respectively. The height of crane post is 3 m and tie rod remains horizontal. Determine the forces produced in the jib and tie rod when a load of 2 kN is suspended at the crane head


1
Expert's answer
2021-10-10T09:32:53-0400

Here it is given vertical post is 3 m high length of jib is 5 m long and tie is 4 m long weight acting is 2 kN since the line of action of all forces acts on a coincident point


"h_r =5 -3 = 2m\\
tan\theta= 3/2\\
\theta =56.31°\\
T =2/\cos56.31°= 3.606\textsf{ kN}"